at what rate of simple interest a principal amount becomes three time in 20 years
Answer by fcabanski(1391)

i=Prt where i is the interest collected, P is the starting principal, r is the interest rate, and t is the time. Since the time is given in years, the interest is considered as yearly interest rate.


Call the starting Principal x, since it is unknown. Add the principal, x, to the accumulated interest, to have three times the original amount. x is the original amount, 3x is three times the original amount, so 2x interest is needed.


2x = x*20*r


2x=20xr
2x/20x=r=1/10 = .1 = 10% interest.